OpenStudy (anonymous):

Determine the measure of the base angles of an isosceles triangle in which the equal sides are 2/3 of the length of the base.

OpenStudy (anonymous):

First you need to know isosceles means two of the triangles sides (and therefore angles) are equal and there is 180 degrees in a triangle. so call the angle of the two equal sides x and x. So we can say x + x + 2/3x = 180 so 2 2/3x = 180 x = 67.5. So the angle we want to find is 2/3 * 67.5 = 45.